OpenStudy (anonymous):

Giselle is going to frame a portrait of the family and place it on the mantle in the family room. The portrait is 10 inches longer than it is tall and will take up a total area of 1344 square inches once it is inside a frame that adds 2 inches to the portrait's width and height. Find the dimensions and area of the unframed portrait.

OpenStudy (aravindg):

\[\huge \text{WELCOME TO OPENSTUDY!}\]

OpenStudy (whpalmer4):

Write the equations that represent the story: Let Lp be the length of the portrait, and Wp be the width Let Lf be the length of the frame, and Wf be the width \[L_f * W_f = 1344\]\[L_p = W_p + 10\]\[L_f=L_p+2\]\[W_f=W_p+2\] Now combine them to get an equation in terms of Lf and Lp which you should be able to solve for one, then use your solution to find the other.